OpenStudy (anonymous):

Find the perimeter of a rectangle with l = 3.2 m and w = 0.5 m. A. 1.6 m B. 3.2 m C. 3.7 m D. 7.4 m

jhonyy9 (jhonyy9):

hope you know perimeter = 2l +2w ok ?

jhonyy9 (jhonyy9):

2l = 6,4 2w = 1 2l +2w = 6,4 + 1 =7,4 hope is understandably